OpenStudy (anonymous):

Write a direct variation that relates x to y. Then show both equations and solution. If y=8 and x=4, find y when x=5

OpenStudy (anonymous):

since y varies directly with x therefore \[y \alpha x \rightarrow y =kx\] where k is any constant now at y=8 and x=4 we need to find k so \[8=k \times 4\rightarrow k = \frac{8}{4} \rightarrow k=2\] Now substituting k =2 in y =kx we find y=2x Now at x=5 we are required to find the value of y so y= 2*5=10 i.e. y =10 Thus at x=5, y=10

OpenStudy (anonymous):

ty sir/mam